OpenStudy (anonymous):

I need help! I am stuck and I do not know how to do this... can someone please help me work through this...This is an essay question.. Idk where to begin... Given the function f(x) = 3(x − 3)2 + 2, indicate the shifts that will affect the location of the vertex and explain what effect they will have. Use complete sentences. f(x+4) f(x) + 4 f(4x) 4•f(x)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

\[y = f(x)\]We can make that puppy sit up and do tricks :-) What happens if we add or subtract something from the result of a function?

OpenStudy (anonymous):

I am really not sure.

OpenStudy (whpalmer4):

If we have \(f(x) = x\), simple case, a straight line going through the origin and up and to the right with a slope of 1.

OpenStudy (anonymous):

so if you add or subtract something from the final result of a function it determines if it moves up or down?

OpenStudy (whpalmer4):

Now let's add 1 to the function: \[y = f(x) +1 = x+1\]What does our new graph look like? What is the value of \(y\) at \(x = 0\)? \(x=1\), \(x=-3\)?

OpenStudy (whpalmer4):

Yes, adding a positive value to the result translates the graph in the direction of positive \(y\), and adding a negative value translates in the direction of negative \(y\). Very good!

OpenStudy (whpalmer4):

Now here's a trickier case: how do we shift right or left?

OpenStudy (anonymous):

Uhhh now that I am really not sure of.... I do not even have a guess.

OpenStudy (anonymous):

and is that x=0?

OpenStudy (whpalmer4):

Let's make a table of values: x y=f(x)=x y = f(x+1) = x+1 ------------------------------- -3 -3 -2 -2 -2 -1 -1 -1 0 0 0 1 1 1 2 2 2 3 3 3 4

OpenStudy (whpalmer4):

Imagine if you plotted both of the right hand columns? what would you get?

OpenStudy (anonymous):

What do you mean what would I get? Would I have a parabola?

OpenStudy (whpalmer4):

okay, here are the two lists of points: (-3,-3), (-2,-2), (-1, -1), (0,0), (1,1), (2,2), (3,3) that's the first line (-3,-2), (-2, -1), (-1, 0), (0, 1), (1,2), (2,3), (3,4) that's the second line

OpenStudy (whpalmer4):

aren't the y values in the y = f(x+1) column identical to the values in the y = f(x) column, except shifted up or down by one row?

OpenStudy (anonymous):

ohhh yess. I see that.

OpenStudy (whpalmer4):

Right. so the effect of adding or subtracting from the ARGUMENT of the function is to translate the graph to the left or the right.

OpenStudy (anonymous):

Okay that makes sens

OpenStudy (whpalmer4):

if we subtract a positive number from the argument, instead of getting a graph feature that normally takes place at x = 0, we'll end up with the graph feature that takes place at x = <number we subtracted>.

OpenStudy (whpalmer4):

and if we add a positive number, instead of getting a graph feature that normally takes place at x = 0, we'll end up with the graph feature that takes place at x = -<number we added>. refer back to my 3 column table and convince yourself that I speak truth :-)

OpenStudy (whpalmer4):

it's easy and logical for the up/down translation — adding makes it go up, subtracting makes it go down. You would think that the natural order of things would have adding to the argument shifting the graph to the right, but in fact, it shifts to the left!

OpenStudy (anonymous):

Okay (:

OpenStudy (anonymous):

This makes much more sense now (:

OpenStudy (whpalmer4):

So, now we can translate our graph right/left, and up/down. What about combination? What if we want to shift \(y = f(x)\) 2 units up, and 3 to the left? What would the new function be?

OpenStudy (anonymous):

would I have f(x+2)-3 ?

OpenStudy (whpalmer4):

well, the right pieces, but not quite the right arrangement. 2 units up, 3 to the left means we have to add 2 to the result of the function, and add 3 to the argument. Here's a plot of 3 parabolas. In blue, we have \(y = f(x) = x^2\). In purple, we have \( y = f(x+3)+2 = (x+3)^2+2\) and in olive, we have \(y = f(x+2)-3 = (x+2)^2-3\)

OpenStudy (anonymous):

oh okay (:

OpenStudy (anonymous):

Hey, I am sorry to leave while you are in the middle of helping me... because it is rude. But my mom just called me to go do something, so I will be back.

OpenStudy (anonymous):

Again I am sorry.

OpenStudy (whpalmer4):

No problem! I may continue on without you :)

OpenStudy (anonymous):

Okay I am back and ready to learn! lol

OpenStudy (whpalmer4):

Oh, words to warm my heart :-)

OpenStudy (anonymous):

hahaha xp

OpenStudy (whpalmer4):

Hey, I'm not kidding, you know! So looking back at your answer, do you see where you went wrong?

OpenStudy (anonymous):

Not really.... I honestly am not understanding this... I am confused. I swear I am trying to understand... but my brain does not want to.

OpenStudy (whpalmer4):

Okay. Let's recap: adding to the result of the function translates up/down: \(y = f(x)+a\) looks identical to \(y = f(x)\), except it appears \(a\) units higher on the y-axis. If \(a\) is negative, then it will be a negative amount higher (read: lower :-)

OpenStudy (whpalmer4):

That's the easiest case to understand, so let's make sure it is crystal-clear.

OpenStudy (anonymous):

Okay so if you add to a function it translates upward and if you subtract it translates down?

OpenStudy (whpalmer4):

yes, assuming we are adding or subtracting positive quantities.

OpenStudy (whpalmer4):

if we are adding or subtracting negative quantities, it moves in the opposite direction... \[y = f(x)+(-1) = f(x) - 1\]right?

OpenStudy (anonymous):

Right.... I think I understand lol

OpenStudy (whpalmer4):

sometimes I'll ask questions which are really quite simple, like that one :-) Don't be fooled into thinking that you're missing something more complicated if it seems too simple to be true :-)

OpenStudy (whpalmer4):

\[y = f(x)+a\]translates the graph along the \(y\) axis by \(a\), with \(a>0\) moving the graph in the direction of increasing \(y\) values, and \(a<0\) moving the graph in the direction of decreasing \(y\) values. What happens if \(a=0\)?

OpenStudy (anonymous):

Does it stay the same?

OpenStudy (whpalmer4):

Very good. That might qualify as one of those questions I was talking about :-)

OpenStudy (anonymous):

Okay haha (: I understand this a little better.

OpenStudy (whpalmer4):

Now, let's wrestle with the right-left business again. \[y=f(x+a)\]translates the graph along the \(x\) axis by \(a\), with \(a>0\) moving the graph to the LEFT and \(a<0\) moving the graph to the RIGHT

OpenStudy (anonymous):

Okay (:

OpenStudy (whpalmer4):

Suppose the graph has some interesting feature at \(x=3\). Maybe it touches the \(x\) axis there, or has a vertex, or whatever. Something readily recognizable. If we add \(a\) to the argument before we put it in the \(f(x)\) machinery (in other words, we calculate \(f(x+a)\) instead of \(f(x)\), that thing we previously saw at \(x=3\) is now going to appear over at \(x = 3-a\), agreed?

OpenStudy (whpalmer4):

if \(a=0\), the graph stays exactly where it started, of course. But if we added \(a\), to get the same result out of the machine, we need to put in \(x-a\), so that we have \(f((x-a)+a) = f(x - a + a) = f(x)\)

OpenStudy (anonymous):

Yes. That makes sense.

OpenStudy (whpalmer4):

I'm going to repost your question here so I don't have to scroll quite so far back: Given the function\( f(x) = 3(x − 3)^2 + 2\), indicate the shifts that will affect the location of the vertex and explain what effect they will have. Use complete sentences. \[f(x+4)\]\[ f(x) + 4\]\[ f(4x)\]\[ 4•f(x)\]

OpenStudy (anonymous):

Okay (:

OpenStudy (whpalmer4):

Okay, we've covered the first two types of transformations. Now let's do the last two types mentioned here: \[y = f(ax)\]\[y = af(x)\]

OpenStudy (anonymous):

Alright (:

OpenStudy (whpalmer4):

Now, before we do it, any guesses as to what multiplying either the argument or the result is going to do to the graph? We've already got shifting in each of the compass directions covered, so it's not that. What else might we do?

OpenStudy (anonymous):

ummm do you like double shift or something lol? I have no idea lol xp

OpenStudy (anonymous):

Do you stretch the graph???? possibly??

OpenStudy (whpalmer4):

yes, you stretch or compress it, very good!!!

OpenStudy (anonymous):

Haha... I feel smart xp

OpenStudy (whpalmer4):

Let's do some experiments. I like experiments :-) Let's say we have a function \(y = \sqrt{1-x^2}\) which looks a bit scary, but is just the top half of a circle with radius 1, centered at the origin \((0,0)\)

OpenStudy (anonymous):

Okay (:

OpenStudy (whpalmer4):

What do you think will happen if instead of plotting \(y = f(x) = \sqrt{1-x^2}\) I instead plot \[y = 2*f(x) = 2*\sqrt{1-x^2}\]?

OpenStudy (anonymous):

Will the graph be stretched by 2?

OpenStudy (anonymous):

stretched up*

OpenStudy (anonymous):

Oh cool so I was right! xp

OpenStudy (whpalmer4):

Here's one where I multiplied by 1/2 instead of 2

OpenStudy (whpalmer4):

let's see if I can trick you. What will \[y = -2f(x)\] look like?

OpenStudy (anonymous):

wait so it got compressed when it was multiplied by 1/2

OpenStudy (whpalmer4):

Yep!

OpenStudy (anonymous):

Oh okay (: Why?

OpenStudy (whpalmer4):

well, everything was half as tall, right?

OpenStudy (whpalmer4):

it's a linear stretching or compressing along the y-axis

OpenStudy (anonymous):

Yes. Okay that makes sense (: ..... anyway, if it is multiplied by -2 it would it be compressed by 2?

OpenStudy (whpalmer4):

if |a|>1, it's a stretch, and |a|<1 a compression

OpenStudy (whpalmer4):

ah, you fell into the trap :)

OpenStudy (whpalmer4):

multiplying by -2 stretches it by a factor of 2, and inverts it!

OpenStudy (anonymous):

Oh wow xp this is complicated lol

OpenStudy (whpalmer4):

I probably wouldn't have fooled you if I had just shown you \[y = -1*f(x) = -f(x)\]first. But where's the fun for me in that? :-)

OpenStudy (whpalmer4):

Changing the sign of the result just flips the graph over the x-axis. What was up, now is down :-)

OpenStudy (anonymous):

hahaha xp At least one of us finds math fun.... just out of curiosity are you a teacher or something?

OpenStudy (whpalmer4):

Nah, I'm an engineer.

OpenStudy (anonymous):

Oh haha well that explains why you are good at math then

OpenStudy (whpalmer4):

I'm on a quixotic quest to show people who think math is hard/they can't do it that it really doesn't have to be, at least for the stuff you learn in junior and senior high school. It's just a matter of getting the right explanation, practice, and careful work. It's very rewarding to get someone to go from "I hate math" to "OMG, I can do this, it's so easy!" :-)

OpenStudy (anonymous):

I see, I see xp .... Well I used to like math until my freshman year of High school when I got a really terrible teacher... and ever since math and I have not been friends.

OpenStudy (whpalmer4):

The problem for many is getting the explanation that works for them. The poor math teacher has to teach what, 20-40 kids at once, has a limited amount of time, and probably has a sucky work environment, too. I get to pick individual students who seem interested in learning, I can take multiple shots at explaining something until I find one that clicks, and I have a really comfy chair in front of a big computer, and the kitchen is just a few feet away :-)

OpenStudy (whpalmer4):

So, do you see why changing the sign inverts the graph, on top of whatever else we might be doing to molest that poor, innocent function?

OpenStudy (anonymous):

No I do not really get why it is inverted

OpenStudy (whpalmer4):

Well, let's look at the table we didn't make: x y -y -3 0 0 -2 1 -1 -1 2 -2 0 3 -3 1 2 -2 2 1 -1 3 0 0 If you plotted the y column, you get something that looks like the upper half of a circle. If you plot the -y column, you get the same thing, except where the top one went up 3, this one goes down 3, and so on. Isn't that how you would describe inversion?

OpenStudy (anonymous):

Ohhh yes... Okay I understand (:

OpenStudy (whpalmer4):

Right. So now we've handled stretching/compressing in the up/down direction, we still need to do that on the right-left direction. I bet you can tell me how...

OpenStudy (anonymous):

Does that involve stretching and compressing to the right and left?? Possibly?

OpenStudy (whpalmer4):

yes. but what might we do to the formula? here's a hint: translate up/down: add to the result stretch/compress up/down: multiply with the result translate left/right: add to the argument stretch/compress right/left: ???